OpenStudy (anonymous):

whats the value of y? 4y+15=6y-45

OpenStudy (mathstudent55):

To solve for y, you need to isolate y. That means, subtract 6y from both sides to have all y-terms on the left side. Also, you need to subtract 15 from both sides to bring all numbers to the right side.

OpenStudy (mathstudent55):

Start by doing the two steps above. What do you get?

OpenStudy (anonymous):

subtract 15 from both sides and 6y from both sides to get -2y=-60 then divide by negative 2.

OpenStudy (anonymous):

-30 ?

OpenStudy (mathstudent55):

Negative divided by negative is positive.

OpenStudy (anonymous):

no, because a negative divided by a negative is a positive.

OpenStudy (anonymous):

so a negative 60 divided by a negative 2 should be what?

OpenStudy (anonymous):

30

OpenStudy (anonymous):

correct

OpenStudy (mathstudent55):

\(4y+15=6y-45\) Subtract 6y from both sides: \(-2y + 15 = -45\) Subtract 15 from both sides: \(-2y = -60\) Divide both sides by -2: \(\dfrac{-2y}{-2} = \dfrac{-60}{-2} \) \(y = 30\)